# On Continuous Optimization for Constraint Satisfaction Problems

### v1

@@ Appendix A Technical Proof
where the second equation holds due to Definition 3. By substituting Eq. (8) and Eq. (9) into Eq. (A), we can have
@@
Proof of Theorem 3
@@
Proof of Corollary 1
$=\mathop{\mathbb{P}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)=1]$
$=\texttt{COP}_{c}(P).$
Hence,
$\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[\mathcal{F}(X)]$ $=\sum_{c\in{}C}\mathop{\mathbb{E}}_{X\sim\mathcal{S}_{P}}[f_{c}(X)]$
$=\sum_{c\in{}C}\texttt{COP}_{c}(P)$
$=\mathcal{C}(P)$
Proof of Lemma 1
We prove by induction on the MDD $G(V,E)$. Let ${P}\in\tilde{D}$.
  • Inductive step: For each non-root node $v$ of $V$, let $par_{i}(v)$ be the set of parent nodes of $v$ with an edge labeled by $i$. After Algorithm 2 terminate, we have
    $m_{td}[v]=\sum_{u\in par_{i}(v)}{p}_{id(u),i}\cdot m_{td}[u],$
    $m_{td}[v]$ $=\sum_{u\in par_{i}(v)}\mathbb{P}[b_{id(u)}=i]\cdot\mathbb{P}[u|G,{P}]$
    $=\mathbb{P}[v|G,{P}]$
Proof of Lemma 2
We prove this by structural induction on MDD $G(V,E)$. Let ${P}\in\tilde{D}$, we have
    $m_{bu}[v]=\sum_{i\in\mathrm{dom}(x_{id(v)})}{p}_{id(v),i}\cdot m_{bu}[v.i].$
    By the inductive hypothesis, we have
    $m_{bu}[v]$ $=\sum_{i=0}^{k}{p}_{id(v),i}\cdot\texttt{COP}_{c,v.i}({P})$
    $=\sum_{i=0}^{k}\mathbb{P}[B_{id(v)}=i]\cdot\texttt{COP}_{c,v.i}({P})$
    $=\texttt{COP}_{c,v}({P})$
    where $\texttt{COP}_{c,v}$ denotes the circuit-output probability of subgraph of $G$ rooted at node $v$.
Proof of Theorem 4
Given an MDD $G(V,E)$, expand all terminal nodes such that each can only have one parent node. The expanded MDD is called $\mathcal{G}$, which has a true terminal set $T$ and a false terminal set $F$. We have the following by repeating the proof of Lemma 1 and 2 on $\mathcal{G}$.
$\texttt{COP}_{c}({P})=\sum_{v\in{}T}\mathbb{P}[v|{P},\mathcal{G}]=\sum_{v\in{}T}m_{td}^{\prime}(v)$
For any non-terminal node $v\notin(T\cup F)$,
$m_{td}[v]$ $=m_{td}^{\prime}[v],$
$m_{bu}[v]$ $=m_{bu}^{\prime}[v].$
For any true terminal node $v\in{T}$ and false terminal node $u\in{F}$,
$m_{bu}^{\prime}[v]$ $=1,$
$m_{bu}^{\prime}[u]$ $=0.$
Now, we are ready to prove Theorem 4.
$\texttt{COP}_{c}({P})$ $=\mathbb{P}[f_{c}=1|{P},\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}\mathbb{P}[v|{P},\mathcal{G}]\cdot\mathbb{P}[f_{c}=1|v,{P},\mathcal{G}]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}^{\prime}[v]\cdot m_{bu}^{\prime}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v]$
$=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot\sum_{j=0}^{k}p_{i,j}m_{bu}[v.j]$
Hence we have
$\frac{\partial\texttt{COP}_{c}(P)}{\partial p_{i,j}}=\sum_{v\in\{u:id(u)=i\}}m_{td}[v]\cdot m_{bu}[v.j]$
Proof of Lemma 3
To prove Lemma 3, we first establish three propositions that collectively support the main result.
Proposition 1 (Lipschitz).
Let $\mathcal{C}:\tilde{D}\to[0,|C|]$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. For every point $y,z\in\tilde{D}$, we have
$|\mathcal{C}(y)-\mathcal{C}(z)|$ $\leq|C|N^{\frac{1}{2}}\|y-z\|,$
i.e., $\mathcal{C}$ is $|C|N^{\frac{1}{2}}$-Lipschitz continuous, and $L=|C|N^{\frac{1}{2}}$ is the Lipshitz constant. As a consequence of the Lipschitz continuity, we have
@@
$|\mathcal{C}(y)-\mathcal{C}(z)-\left<\nabla\mathcal{C}(z),y-z\right>|\leq\frac{L}{2}\|y-z\|^{2}.$ (10)
Proof of Proposition 1.
By the triangle inequality,
$|\mathcal{C}(y)-\mathcal{C}(z)|$
$\leq$ $|\mathcal{C}(y_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})|$
$+$ $|\mathcal{C}(z_{1,1},y_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},z_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})|+\cdots$
$+$ $|\mathcal{C}(z_{1,1},z_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{C}(z_{1,1},z_{1,2},\cdots,z_{n,|\mathrm{dom}(x_{n})|})|$
$=$ $|(y_{1,1}-z_{1,1})\partial_{1,1}\mathcal{C}(y_{1,2},\cdots,y_{n,|\mathrm{dom}(y_{i})|})|+\cdots+|(y_{n,|\mathrm{dom}(x_{n})|}-z_{n,|\mathrm{dom}(x_{n})|})\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{C}(z_{1,1},\cdots,z_{n,|\mathrm{dom}(x_{n})|-1})|$
Due to Theorem 4, $|\partial_{i,j}\mathcal{C}|\leq|C|$ for all $i\in[n]$ and $j\in[|\mathrm{dom}(x_{i})|]$, we have
$|(y_{1,1}-z_{1,1})\partial_{1,1}\mathcal{C}(y_{1,2},\cdots,y_{n,|\mathrm{dom}(y_{i})|})|+\cdots+|(y_{n,|\mathrm{dom}(x_{n})|}-z_{n,|\mathrm{dom}(x_{n})|})\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{C}(z_{1,1},\cdots,z_{n,|\mathrm{dom}(x_{n})|-1})|$
$\leq$ $|C|\left(\sum_{i\in[n]}\sum_{j\in[|\mathrm{dom}(x_{i})|]}|y_{i,j}-z_{i,j}|\right)$
$\leq$ $|C|N^{\frac{1}{2}}\|y-z\|$
The proof of Eq. (10) is given by representing $\mathcal{C}(y)-\mathcal{C}(z)$ as an integral.
@@
$\leq$ $\int_{0}^{1}\|\nabla\mathcal{C}(z+t(y-z))-\nabla\mathcal{C}(z)\|\cdot\|y-z\|dt$
$\leq$ $\int_{0}^{1}L\cdot t\cdot\|y-z\|^{2}dt$
$=$ $\frac{L}{2}\|y-z\|^{2}dt$
The Lipschitz continuity bound the smoothness of $\mathcal{C}$. By performing projected gradient descent, $\mathcal{C}$ will converge to the boundary of the simplices. For each projection step, we have the following Proposition.
Proposition 2 (Projection Point (Nesterov 2013)).
For $y,z\in\tilde{D}$, we define the distance function be
$h(y)=\|y-(z+\eta\nabla\mathcal{C}(z))\|^{2},$
where $\eta>0$ is the step size. Let $y^{*}=\operatorname*{arg\,min}_{y\in\tilde{D}}h(y)$ be the Euclidean projection of the ascent step $z+\eta\nabla\mathcal{C}(z)$ onto $\tilde{D}$. Then the following optimality condition holds:
$\left<\nabla h(y^{*}),y-y^{*}\right>\geq 0.$
In particular, the projection point satisfies $y^{*}=z+\eta{}g(z)$, where $g(z)$ denotes the projected gradient mapping in the ascent direction. Hence, we have the following Proposition.
Proposition 3 (Gradient Mapping).
Given a function $\mathcal{C}$ defined on $\tilde{D}$, for every $y\in\tilde{D}$, we have
$\left<\nabla\mathcal{C}(y),g(y)\right>\geq\|g(y)\|^{2},$
where $g(y)$ is the gradient mapping.
Proof of Proposition 3.
Given any point ${z}\in{\tilde{D}}$, the projection point $y^{*}={z}+\eta{}g({z})$. By substituting $y^{*}$ with ${z}+\eta{}g({z})$ in Proposition 2, for all $y\in\tilde{D}$ we have
$\left<\nabla{}h({z}+\eta{}g({z})),y-({z}+\eta{}g({z}))\right>\geq 0.$
Let y=zy={z}, the above inequality becomes
$\left<\nabla{}h(y+\eta{}g(y)),-\eta{}g(y))\right$> $\geq 0$
$\left<2((y+\eta{}g(y))-(y+\eta{}\nabla\mathcal{C}(y))),-\eta{}g(y)\right$> $\geq 0$
$\left<2\eta{}(g(y)-\nabla\mathcal{C}(y)),-\eta{}g(y)\right$> $\geq 0$
$\left<\eta^{2}(\nabla\mathcal{C}(y)-g(y)),g(y)\right$> $\geq 0.$
For any $\eta>0$, we have $\left<(\nabla\mathcal{C}(y)-g(y)),g(y)\right>\geq 0$
By combining the propositions, we are now ready to quantitatively measure the progress made in every gradient step.
Proof of Lemma 3.
Eq. 10 can be rewritten as
$\mathcal{C}(y)\geq\mathcal{C}(z)+\left<\nabla\mathcal{C}(z),y-z\right>-\frac{L}{2}\|y-z\|^{2}$
Let $y=z+\eta{}g(z)$, the above inequality becomes
$\mathcal{C}(y)\geq\mathcal{C}(z)+\eta\left<\nabla\mathcal{C}(z),g(z)\right>-\frac{\eta^{2}\cdot L}{2}\|g(z)\|^{2}.$
Applying Proposition 3 and let $\eta\leq{}\frac{1}{L}$, we then have
$\mathcal{C}(y)$ $\geq\mathcal{C}(z)+\eta\|g(z)\|^{2}-\frac{\eta}{2}\|g(z)\|^{2}$
$=\mathcal{C}(z)+\frac{\eta}{2}\|g(z)\|^{2}.$
Proof of Theorem 5
From Lemma 3, we can assume $\eta=\frac{1}{L}$. Consider a sequence of points ${P}_{t}\in\tilde{D}$ such that ${P}_{t+1}={P}_{t}+\eta g({P}_{t})$. Due to Lemma 3 we have $\mathcal{C}({P}_{(t+1)})-\mathcal{C}({P}_{t})\geq\frac{\eta}{2}\|g({P}_{t})\|^{2}$. Given any initial point $a_{(0)}$, it will converge to ${P}^{*}$. After $T$ gradient step, we have
$\mathcal{C}({P}^{*})-\mathcal{C}({P}_{0})$ $\geq\frac{1}{2L}\sum_{t=0}^{T}\|g(a_{(t)})\|^{2}$
$\geq\frac{T+1}{2L}\min_{t}\|g(a_{(t)})\|^{2},$
@@
$\min_{t}\|g(a_{(t)})\|$ $\leq\sqrt{\frac{2L}{T+1}}\sqrt{\mathcal{C}({P}^{*})-\mathcal{C}({P}_{0})}$
@@
To achieve a point ${P}_{T}$ such that $\|g({P}_{t})\|\leq\epsilon$, we require $\sqrt{\frac{2|C|L}{T+1}}\leq\epsilon$, which implies $T\sim\mathcal{O}\left(\frac{|C|L}{\epsilon^{2}}\right)$.
Proof of Lemma 4 and Theorem 6
Since $\mathcal{C}$ is a multilinear function, it can be decomposed as
$\mathcal{C}(A)=\sum_{j\in\mathrm{dom}(x_{1})}a_{1,j}{}g_{j}(A_{\backslash{}1})+h(A_{\backslash{}1}),$
where $g_{j}$’s and $h$ are multilinear polynomials not depending on $j$. Let
$\eta q^{+}_{i}=\begin{cases}{{proj}_{\tilde{D}_{i}}(a_{1}+\eta\nabla\mathcal{C}(a_{1}))-a_{1}},&i=1,\\ \vec{0},&otherwise,\end{cases}$
$\mathcal{C}(A+\eta{}q^{+})-\mathcal{C}(A)$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\eta q^{+}_{1,j}\cdot g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left({proj}_{\tilde{D}}(a_{1}+\eta\nabla\mathcal{C}(a_{1}))_{j}-a_{1,j}\right)\cdot g_{j}(A_{\backslash{}1}).$
If ${proj}_{\Delta_{1}}(a_{1}+\eta{}q_{1}^{+})$ is on $\Delta\backslash{\delta\Delta}$, we have
$\frac{1}{\eta}\left(\mathcal{C}(A+\eta{}q^{+})-\mathcal{C}(A)\right)=$ $\frac{1}{\eta}\sum_{j\in\mathrm{dom}(x_{1})}\left(\eta g_{j}(A_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}\left(a_{1,j}+\eta g_{j}(a_{\backslash{}1})\right)-1}{|\mathrm{dom}(x_{1})|}\right)g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left(g_{j}(A_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})}{|\mathrm{dom}(x_{1})|}\right)\cdot g_{j}(A_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})^{2}-\frac{1}{|\mathrm{dom}(x_{1})|}\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})\right)^{2}$
$=$ $\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(A_{\backslash{}1})^{2}\right)\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{1}{|\mathrm{dom}(x_{1})|}\right)-\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{g_{j}(A_{\backslash{}1})}{\sqrt{|\mathrm{dom}(x_{1})|}}\right)^{2}$
$\geq$ $0$
The last inequality is the Cauchy-Schwarz inequality, where the equality holds when $\mathcal{C}(a)$ is a constant.
If ${proj}_{\tilde{D}}(a\pm\eta{}q^{\pm})$ is on $\delta\tilde{D}$}, we can always find an $\epsilon^{\prime}\in(0,\epsilon)$ such that for all $\eta^{\prime}\in(0,\epsilon^{\prime})$, ${proj}_{\tilde{D}}(a\pm\eta^{\prime}q^{\pm})$ is on $\tilde{D}\backslash\delta\tilde{D}$}.
Lemma 4 states that, any interior point $a\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local maximum of $\mathcal{C}(a)$. In contrast, all local maxima must lie on the boundary $\delta\tilde{D}$.

### v2

@@ Appendix A Technical Proof
where the second equation holds due to Definition 3.2. By substituting Eq. (7) and Eq. (8) into Eq. (A.1), we can have
@@
A.2 Proof of Proposition 3.5
By the triangle inequality,
$|\mathcal{F}(P)-\mathcal{F}(P^{\prime})|$
$\leq$ $|\mathcal{F}(P_{1,1},P_{1,2},\cdots,P_{n,|\mathrm{dom}(x_{n})|})-\mathcal{F}(P^{\prime}_{1,1},P_{1,2},\cdots,P_{n,|\mathrm{dom}(x_{n})|})|$
$+$ $|\mathcal{F}(P^{\prime}_{1,1},P_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{F}(P^{\prime}_{1,1},P^{\prime}_{1,2},\cdots,P_{n,|\mathrm{dom}(x_{n})|})|+\cdots$
$+$ $|\mathcal{F}(P^{\prime}_{1,1},P^{\prime}_{1,2},\cdots,y_{n,|\mathrm{dom}(x_{n})|})-\mathcal{F}(P^{\prime}_{1,1},P^{\prime}_{1,2},\cdots,P^{\prime}_{n,|\mathrm{dom}(x_{n})|})|$
$=$ $|(P_{1,1}-P^{\prime}_{1,1})\cdot\partial_{1,1}\mathcal{F}(P^{\prime}_{1,2},\cdots,P^{\prime}_{n,|\mathrm{dom}(x_{n})|})|+\cdots$
$+$ $|(P_{n,|\mathrm{dom}(x_{n})|}-P^{\prime}_{n,|\mathrm{dom}(x_{n})|})\cdot\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{F}(P^{\prime}_{1,1},\cdots,P^{\prime}_{n,|\mathrm{dom}(x_{n})|-1})|$
Since $\mathcal{F}(P)$ is a multilinear function, $|\partial_{i,j}\mathcal{F}|\leq|C|$ for all $i\in[n]$ and $j\in[|\mathrm{dom}(x_{i})|]$. Hence we have
$|(P_{1,1}-P^{\prime}_{1,1})\partial_{1,1}\mathcal{F}(P_{1,2},\cdots,P_{n,|\mathrm{dom}(x_{n})|})|+\cdots$
$+$ $|(P_{n,|\mathrm{dom}(x_{n})|}-P^{\prime}_{n,|\mathrm{dom}(x_{n})|})\partial_{n,|\mathrm{dom}(x_{n})|}\mathcal{F}(P^{\prime}_{1,1},\cdots,P^{\prime}_{n,|\mathrm{dom}(x_{n})|-1})|$
$\leq$ $|C|\left(\sum_{i\in[n]}\sum_{j\in[|\mathrm{dom}(x_{i})|]}|P_{i,j}-P^{\prime}_{i,j}|\right)$
$\leq$ $|C|\sqrt{N}\|P-P^{\prime}\|$
A.3 Proof of Theorem 3.6
@@
A.4 Proof of Lemma 3.7
Since $\mathcal{F}$ is a multilinear function, it can be decomposed as
$\mathcal{F}(P)=\sum_{j\in\mathrm{dom}(x_{1})}p_{1,j}{}g_{j}(P_{\backslash{}1})+h(P_{\backslash{}1}),$
where $g_{j}$’s and $h$ are multilinear polynomials not depending on $j$. Let
$\eta q_{i}=\begin{cases}{{proj}_{\tilde{D}_{i}}(p_{1}-\eta\nabla\mathcal{F}(p_{1}))-p_{1}},&i=1,\\ \vec{0},&otherwise,\end{cases}$
$\mathcal{F}(P-\eta{}V)-\mathcal{F}(P)$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\eta v_{1,j}\cdot g_{j}(P_{\backslash{}1})$
@@
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left({proj}_{\tilde{D}}(p_{1}-\eta\nabla\mathcal{F}(p_{1}))_{j}-p_{1,j}\right)\cdot g_{j}(P_{\backslash{}1}).$
If ${proj}_{\Delta_{1}}(p_{1}-\eta{}q_{1})$ is on $\Delta\backslash{\delta\Delta}$, we have
$\frac{1}{\eta}\left(\mathcal{F}(P-\eta{}V)-\mathcal{F}(P)\right)$
$=$ $\frac{1}{\eta}\sum_{j\in\mathrm{dom}(x_{1})}\left(\eta g_{j}(P_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}\left(p_{1,j}+\eta g_{j}(P_{\backslash{}1})\right)-1}{|\mathrm{dom}(x_{1})|}\right)g_{j}(P_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}\left(g_{j}(P_{\backslash{}1})-\frac{\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(P_{\backslash{}1})}{|\mathrm{dom}(x_{1})|}\right)\cdot g_{j}(P_{\backslash{}1})$
$=$ $\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(P_{\backslash{}1})^{2}-\frac{1}{|\mathrm{dom}(x_{1})|}\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(P_{\backslash{}1})\right)^{2}$
$=$ $\left(\sum_{j\in\mathrm{dom}(x_{1})}g_{j}(P_{\backslash{}1})^{2}\right)\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{1}{|\mathrm{dom}(x_{1})|}\right)-\left(\sum_{j\in\mathrm{dom}(x_{1})}\frac{g_{j}(P_{\backslash{}1})}{\sqrt{|\mathrm{dom}(x_{1})|}}\right)^{2}$
$\geq$ $0$
The last inequality is the Cauchy-Schwarz inequality, where the equality holds when $\mathcal{F}$ is a constant.
A.5 Proof of Theorem 3.8
From Lemma 3.7 we can observe that, if ${proj}_{\tilde{D}}(P-\eta{}V)$ is on $\delta\tilde{D}$}, we can always find an $\epsilon^{\prime}\in(0,\epsilon)$ such that for all $\eta^{\prime}\in(0,\epsilon^{\prime})$, ${proj}_{\tilde{D}}(P-\eta^{\prime}V)$ is on $\tilde{D}\backslash\delta\tilde{D}$}.
Lemma 3.7 states that, any interior point $a\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local maximum of $\mathcal{C}(a)$. In contrast, all local maxima must lie on the boundary $\delta\tilde{D}$. Hence, Lemma 3.7 directly results in Theorem 3.8.
A.6 Proof of Theorem 3.9
To prove Theorem 3.9, we first need to establish Lemma A.3. The proof of Lemma A.3 relies on three propositions.
Proposition A.1 (Projection Point (Nesterov, 2013)).
For $P,P^{\prime}\in\tilde{D}$, let $h(P)=\|P-(P^{\prime}+\eta\nabla\mathcal{F}(P^{\prime}))\|^{2}$ and $P^{*}=\operatorname*{arg\,min}_{P\in\tilde{D}}h(P)$, we have
$\left<\nabla{}h(P^{*}),P-P^{*}\right>\geq 0.$
Note that the projection point $P^{*}$ is $P^{\prime}-\eta{}g(P^{\prime})$; Hence, we have the following Proposition.
Proposition A.2 (Gradient Mapping (Kyrillidis et al., 2021a)).
Given a function $\mathcal{F}$ defined on a convex set $\Omega$, for every $P\in\Omega$, we have
$\left<\nabla\mathcal{F}(P),g(P)\right>\geq\|g(P)\|^{2},$
where $g(P)$ is the gradient mapping.
Also, the proof relies on the consequence of Proposition 3.5. By representing $\nabla\mathcal{F}(y)-\nabla\mathcal{F}(z)$ as an integral, we have
@@
[…]
$\leq$ $\int_{0}^{1}\|\nabla\mathcal{F}(P^{\prime}+t(P-P^{\prime}))-\nabla\mathcal{F}(P^{\prime})\|\cdot\|P-P^{\prime}\|dt$
$\leq$ $\int_{0}^{1}Lt\|P-P^{\prime}\|^{2}dt$
$=$ $\frac{L}{2}\|P-P^{\prime}\|^{2}dt$ (9)
Lemma A.3.
Let $\mathcal{F}$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. Let $g(\cdot)$ be the projected gradient mapping onto the simplices $\tilde{D}$. Given $P,P^{\prime}\in\tilde{D}$, if the step size satisfies $\eta\leq\frac{1}{L}$, then the projected gradient step $P=P^{\prime}-\eta{}g(P^{\prime})$ satisfies:
$\mathcal{F}(P)-\mathcal{F}(P^{\prime})\geq\frac{\eta}{2}\|g(P^{\prime})\|^{2}$
Proof of Lemma A.3.
Eq. (9) can be rewritten as
$\mathcal{F}(P)\leq\mathcal{C}(P^{\prime})+\left<\nabla\mathcal{F}(P^{\prime}),P-P^{\prime}\right>+\frac{L}{2}\|P-P^{\prime}\|^{2}$
Let $P=P^{\prime}-\eta{}g(P^{\prime})$, the above inequality becomes
$\mathcal{F}(P)\leq\mathcal{F}(P^{\prime})-\eta\left<\nabla\mathcal{F}(P^{\prime}),g(P^{\prime})\right>+\frac{\eta^{2}L}{2}\|g(P^{\prime})\|^{2}.$
Applying Proposition A.2 and let $\eta\leq{}\frac{1}{L}$, we then have
$\mathcal{F}(P)$ $\leq\mathcal{F}(P^{\prime})-\eta\|g(P^{\prime})\|^{2}+\frac{\eta}{2}\|g(P^{\prime})\|^{2}$
$=\mathcal{F}(P^{\prime})-\frac{\eta}{2}\|g(P^{\prime})\|^{2}.$
From Lemma A.3, we can assume $\eta=\frac{1}{L}$. Consider a sequence of points $P_{t}$’s such that $P_{t+1}=P_{t}-\eta g(P_{t})$. Due to Lemma A.3 we have $\mathcal{F}(P_{t+1})-\mathcal{F}(P_{t})\leq-\frac{\eta}{2}\|g(P_{t})\|^{2}$. Given any initial point $P_{0}$, it will converge to $P^{*}$. After $T$ gradient steps, we have
$\mathcal{F}(P^{*})-\mathcal{F}(P_{(0)})$ $\geq\frac{1}{2L}\sum_{t=0}^{T}\|g(P_{t})\|^{2}$
$\geq\frac{T+1}{2L}\min_{t}\|g(P_{t})\|^{2},$
@@
$\min_{t}\|g(P_{t})\|$ $\leq\sqrt{\frac{2L}{T+1}}\sqrt{\mathcal{F}(P^{*})-\mathcal{C}(P_{0})}$
@@
To achieve a point $a_{T}$ such that $\|g(P_{t})\|\leq\epsilon$, we require $\sqrt{\frac{2|C|L}{T+1}}\leq\epsilon$, which implies $T\sim\mathcal{O}\left(\frac{|C|\cdot L}{\epsilon^{2}}\right)$.
A.7 Proof of Theorem 3.10
We refer to the full convergence proof in Appendix B.2 of (Allen-Zhu and Orecchia, 2014). $\rho$ was shown in Proposition 3.5 Here, we justify the expression of $\Theta=\sum_{i=1}^{n}\log|\mathrm{dom}(x_{i})|$ as the Bregman diameter under the negative-entropy potential, i.e., $\psi(P)=\langle P,\log P\rangle$. Hence the quantity $\Theta$ is defined as
$\Theta=\sup_{P}\left(\psi(P)-\psi(P_{0})\right),$
where we choose $P_{0}$ to be the uniform distribution on each probability simplex, since this minimizes $\psi(P_{0})$ and therefore yields the tightest bound. Hence,
$\psi(P)$ $=\sum_{i=1}^{n}\sum_{j=1}^{|\mathrm{dom}(x_{i})|}p_{i,j}\log p_{i,j}$
$\psi(P_{0})$ $=\sum_{i=1}^{n}\sum_{j=1}^{|\mathrm{dom}(x_{i})|}\frac{1}{|\mathrm{dom}(x_{i})|}\log\frac{1}{|\mathrm{dom}(x_{i})|}$
$\Theta$ $=\sup_{P}\left(\psi(P)-\psi(P_{0})\right)$
$=\max_{P}\left(\psi(P)\right)-\psi(P_{0})$
$=\sum_{i=1}^{n}\max_{p_{i}}\left(\sum_{j=1}^{|\mathrm{dom}(x_{i})|}p_{i,j}\log p_{i,j}\right)+\sum_{i=1}^{n}\log|\mathrm{dom}(x_{i})|$
$=\sum_{i=1}^{n}\log|\mathrm{dom}(x_{i})|$